OpenStudy (anonymous):

christine went shopping and bought each of her 7 nephews a gift either a video costing 14.95 or a cd costing 16.88 she spent 80.54 on the gifts how many videos and how many cds did she bye

OpenStudy (here_to_help15):

v+c=9 or v=9-c 14.95v+16.88c=138.41 14.95(9-c)+16.88c=138.41 134.55-14.95c+16.88c=138.41 1.93c=138.41-134.55 1.93c=3.86 c=3.86/1.93 c=2 cds were bought. v=9-2=7 video were bought. Proof: 14.95*7+16.88*2=138.41 104.65+33.76=138.41 138.41=134.41

OpenStudy (here_to_help15):

Your welcome

OpenStudy (anonymous):

that was awesome thanks